0.9 g of solid fuel was burnt in a calorimeter. It was observe that the temperature of 4000 g of water in the calorimeter measured by 3.8°c. The composition of hydrogen in the fuel was 1%. The equivalent weight of water is 385g and the latent heat of steam is 587 Cal/g. Calculate: 1. The high calorific value.

2. Low calorific value.


1
Expert's answer
2021-04-14T06:14:33-0400

1. The high calorific value = (4000×3.8)=15200 cal/g

15200- (0.9×1×587) = 14671.7 cal/g


2. Low calorific value = 385×3.8 = 1463 cal/g

1463-(0.9×1×587)= 934.7cal/g